OpenStudy (anonymous):

Simplify: x^2-9/x^2+5x+6

OpenStudy (anonymous):

\[\frac{x^2-9}{x^2+5x+6}\] \[\frac{(x-3)(x+3)}{(x+3)(x+2)}=\frac{x-3}{x+2}\]